OpenStudy (anonymous):

Please help solving the following logarithm: 8^(x-9) = 19^(3x+2)

OpenStudy (anonymous):

So far, I have log8^(x-9) = log19^(3x+2) (x-9)log8 = (3x+2)log19

OpenStudy (mathteacher1729):

so far so good. :) Next step = distribute log(8) over (x-9) and distribute log(19) to (3x+2).

OpenStudy (anonymous):

ok...got that...xlog8-9log8 = 3xlog19 + 2log19

OpenStudy (anonymous):

i always miss it at this step.

OpenStudy (mathteacher1729):

Hint -- it's best to put parenthesis around log(number). log is a function, and the number is the input or independent variable. :) Ok, so the next step here is to "get all the xs on one side" and "get all the log(number) on the other."

OpenStudy (anonymous):

(3xlog19-xlog8) = (9log8-2log19) ???

OpenStudy (mathteacher1729):

Yes. Only thing, use parenthesis around the log of the number... (3xlog(19)-xlog(8)) = (9log(8)-2log(19)) instead of (3xlog19-xlog8) = (9log8-2log19) This is because log(19), log(8) -- pronounced "Log of 19", "log of 8" etc... are numbers. So they are just the same as... if you like, 2 or 3. In the sense they are numbers or constants. Ok, next step -- factor out the x from the expression on the left. (3xlog(19)-xlog(8)) = ?? (factor out the x)

OpenStudy (anonymous):

x(3log(19) - log(8))

OpenStudy (mathteacher1729):

Yup! Now that is still equal to (9log(8)-2log(19)) So you now have x(3log(19) - log(8)) = (9log(8)-2log(19)) It's important to keep in mind that 3log(19) is just a number, as is log(8). So their difference is just a number, it's sort of like saying... x*4 = (9log(8)-2log(19)) What would you do next?

OpenStudy (anonymous):

diviide

OpenStudy (mathteacher1729):

Yup. :) The final answer will be a big fraction with logs in the numerator and denominator.

OpenStudy (anonymous):

i divided 9log(8)-2log(19) on both sides, is that correct

OpenStudy (anonymous):

i'm confusing myslef all over again

OpenStudy (anonymous):

is the answer x=2log(19)+9log(8)/3log(19)-log8

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

thank you